# Thermoelectric properties of defective scandium nitride nanostructures

### v1

@@ III Results and discussion
where $L$ is the length of the central conductor, $A$ is the NW cross-sectional area, and $MP$ is a model parameter obtained by fitting experimental data. This heuristic parameter accounts for (i) the difference in dimensionality between our nanoscale models and the larger experimental thin-film system, (ii) the effect of a finite mean free path in real materials relative to the ballistic transport assumed in our simulations, (iii) the approximate correction of the band gap we employed, and (iv) the compensation for the partially arbitrary choice of an auxiliary function in the model (which will be described later in the text).
@@
We relate the electronic conductances of our NW models to experimental electric conductivity data available in the literature. In Figure 12, panel b, it can be observed that the BT 155v systems exhibit the highest conductivity among the systems we have examined. We associate this model with an hypothetical real ScN system characterized by a strong predominance of N-site vacancies occurring in contiguous positions. The experimental ScN sample from Ref. 24 exhibits some of the highest electric conductivity values ever reported for ScN thin films.
@@
We assume that a strong presence of structural vacancies at N-sites, associated in contiguous positions, is at the origin of the high experimental conductivity. It should be noted that the association of vacancies at N-sites in contiguous positions is not thermodynamically favored. We treat surface defects with the idea to describe the deposition process, and in order to quantify the energy involved, we performed calculations of the formation enthalpies of isolated surface vacancies at N-sites in a ScN slab structure, as well as the formation of associated defects  [^1] We calculated the formation energies of associated defects in 6-atomic-layer slab structures by means of DFT as described in section II.3. The periodic unit used for these calculations had in-plane dimensions of 3×3 conventional simple cubic cells. Atomic layer 1 was kept fixed during the relaxation calculations, and defects were inserted in layers 4 and 6 at the in-plane center of the periodic unit, in axial positions perpendicular to the slab.. The formation enthalpy of two contiguous vacancies results higher by 8.92% compared to that of two isolated vacancies (5.44 eV per single isolated vacancy). However, the formation of a single vacancy at N-site is itself thermodynamically unfavorable, and the first studies on the stability of nitrogen vacancy defects [53] argued that these are not energetically favored and were therefore expected not to be significantly present in ScN lattices. More recent experimental reports have clearly shown, instead, that nitrogen vacancies are prevalent, especially under scandium-rich growth conditions [56, 57]. Although these conditions are not easily achieved with sputtering techniques, this suggests the existence of complex mechanisms for the stabilization of nitrogen vacancies that are not yet fully understood. For instance, the formation of void regions in the bulk structure of ScN has been shown to facilitate a higher concentration of n-type impurities in their vicinity [58], and other theoretical studies have demonstrated a negative formation energy for surface N-site vacancies under Sc-rich conditions [59].
@@
These reasons moved to our assumption and therefore we derived a single model parameter $MP$, which we used for all the systems presented in Figure 12, by fitting the BT 155v conductivity to the experimental data reported in Ref. 24, corresponding to the sample referred to in that work as ScN (here we denote it as Sample I). By intersecting a model-internal theoretical auxiliary function, which we report in Figure 12, panel b (155v aux.), we derived the chemical potentials of that sample at various temperatures from the experimental Seebeck coefficients reported in the same work [24]. As auxiliary function, we used the theoretical Seebeck function of each NW model itself, calculated at a temperature of 450 K (this choice is partially arbitrary and is compensated by the derivation of a suitable $MP$, as previously discussed). We obtained the best possible fit of the model by associating the auxiliary function of the 155 model with the experimental sample labeled ScN-T in Ref. 24 (here denoted Sample II), the auxiliary function of the BT 155v model with the experimental sample labeled ScN in Ref. 24 (Sample I, used to fit $MP$), and the auxiliary function of BT 155O+gl with the experimental sample examined in Ref. 20 (here denoted Sample III). In Figure 12, we report these three auxiliary functions and the calculated conductivities comparing with experimental values from Refs. 20 and 24.
@@
We chose to compare the two models of defected structures noted before that present the most interesting features when comparing their Seebeck function profiles with that of the pristine 155 model, namely: BT 155v and BT 155O+gl. The compared experimental samples (Refs. 20 and 24), deriving from different experimental procedures, differ widely from each other in their characteristics, presenting a wide variability in Seebeck coefficients, electric conductivities, and their variance with temperature. We notice a good agreement between the experimental conductivity values measured for the ScN-T sample in Ref. 24 (Sample II) and the theoretical functions calculated using the NW model 155O+gl (we mention a slightly lower, yet still good, agreement when using the auxiliary function of the 155O model, not reported in figure).
@@
We can therefore describe, within a coherent framework, an interaction between two independent effects that determine the different thermoelectric behaviors of the ScN and ScN-T samples in Ref. 24 (Sample I and II). The different experimental procedures lead, on one hand, to a reduction in v+v defects in ScN-T compared to ScN. This reduction underlies the strong increase in the absolute value of the Seebeck coefficient observed in Ref. 24 and also contributes to the decrease in electrical conductivity. On the other hand, an additional detrimental effect on electronic transport arises from a greater presence of O+gl defects.
@@
Further complexity is detected when analyzing the experimental sample from Ref. 20 (Sample III). This sample also exhibits high conductivity values, so it is reasonable to expect its experimental values to match the highest conductivity function, namely the one calculated using the 155v model. However, in order to achieve good agreement, a different auxiliary function must be used in this case. The best match is obtained when using an auxiliary function calculated from the BT 155O+gl system. The choice of a different auxiliary function accounts for structural effects that may influence the Seebeck coefficient but have little impact on the conductivity. These considerations suggest that different defected structures can coexist within the same sample, acting independently on the electric conductivity and the Seebeck coefficient, with possible contrasting effects.
@@
At this point, it should be noted that while the 155v model has a physically unique structure determined by a structural optimization calculation, the structure used for the 155O+gl model involves a greater degree of arbitrariness in the way the glide-type defect was constructed. As previously discussed, the thermoelectric properties of the 155O+gl model are highly structure-dependent. From an analytical perspective, the shape of its auxiliary function is well suited to fit the model. However, to achieve better numerical agreement, one would need to use the Seebeck function calculated at 600 K as auxiliary function (instead of the one at 450 K, which we use). We instead choose to preserve the internal consistency of the model by calculating all auxiliary functions in the same way. As a result, this set of points yields a lower agreement between theory and experiment compared to the others. However, the reasons for this outcome are clear, as the 155O+gl defect category inherently includes a greater amount of structural variability.
Figure 9: Panel a: Calculated electronic conductances at 0 K as functions of the chemical potential of (1x)5x5+s ScN NW systems: pure (155); BT defected system with an O atom replacing the central N site (155O): relaxed and unrelaxed; and a BT defected system in which a glide-type defect is contiguous to the central O impurity and perpendicular to NW (155O+gl). Panel b: Calculated Seebeck coefficients as functions of the chemical potential at 400 K for the three systems 155, BT 155O (relaxed) and BT 155O+gl. Vertical dashed lines represent the calculated Fermi levels of the systems.
Figure 11: Comparison of the calculated conductances as functions of the chemical potential for the 3×(3×3)+s ScN NW models: pristine BT 333 and TL model composed by 333vv as central conductor and pristine 333 as leads (333-333vv-333).
In Figure 13, we report the calculated Seebeck functions for selected BT 3×(3×3)+s ScN NW systems. These functions were obtained using the same procedure described above, by rigidly shifting the band energies to reproduce the 0.9 eV experimental gap. It can be observed that this procedure yields numerically reasonable results when compared with experimental values. However, this approximation overestimates the distance between the calculated chemical potential and the peak values of the Seebeck functions. Indeed, reducing the cross-section of the NW models leads to increasingly higher absolute peak values in the calculated Seebeck functions, and similar effects are observed when a rigid shift is applied to the energy bands.
@@
Figure 12: Panel a: Auxiliary functions employed to relate the experimental Seebeck coefficients to theoretical values of chemical potential. The figure reports the functions of pristine and two different BT defected (1x)5x5+s ScN NW structures (v+v and O+gl). These functions are calculated as the Seebeck functions at 450 K. The figure also reports, for a comparison, the Seebeck function calculated at 600 K for the BT 155O+gl structure. Panel b: The lines represent calculated electric conductivities as functions of the chemical potential for the same three model systems, at different temperatures. Color intensities of the lines increase with temperature from 300 to 780 K. The points represent, at their calculated chemical potentials, experimental values of electric conductivities: Sample I: ScN from Ref. 24, Sample II: ScN-T from Ref. 24 and Sample III: ScN from Ref. 20.
@@ IV Conclusions
We wondered what the microscopic origins of the experimental variability in thermoelectric response of ScN thin films might be, hypothesizing that they could arise from imperfections and impurities in the lattice structure of ScN. We approached the problem by classifying the possible structural defects and impurities based on their symmetry and chemical nature, and we systematically analyzed those most likely to occur under the experimental conditions typically used.
From the point of view of defect symmetry, we considered isolated point defects, multiple contiguous point defects, and point defects associated with glide-type defects, while from a chemical perspective we studied the presence of oxygen or structural vacancies replacing nitrogen atoms in the lattice. We used the Landauer approach to determine the effect of each of these different microscopic structures on the emergent thermoelectric properties of the whole material.
We highlight the complexity of the problem, aiming to establish a rational framework to explain the different experimental responses. We identified two notable categories of structures whose presence has a considerable impact on the Seebeck coefficient and electrical conductivity at the macroscopic level and which could, if significantly present in the structure, account for the experimentally observed effects.
Specifically, we refer first to the v-v structures, where two or more structural vacancies replace nitrogen atoms at adjacent conventional cells (i.e., N-site vacancies separated by only one Sc atom), which can significantly reduce the absolute value of the Seebeck coefficient while enhancing the electrical conductivity. Second, we refer to the O-gl structures, where an oxygen atom replaces a nitrogen lattice atom, and a glide-type defect is immediately adjacent, involving a displacement of the atomic layer following the one where the substitution occurs. This second category of structures has potentially opposite effects with respect to the firts one, detrimenting the conductivity while enhanching the Seebeck coefficient.

### v3

@@ III.3 Electronic conductivities
where $L$ is the length of the central conductor, $A$ is the NW cross-sectional area, and $MP$ is a model parameter obtained by fitting experimental data. The Landauer approach, with the inclusion of a finite mean free path for backscattering, has been shown to be formally equivalent to the Boltzmann transport equation within the relaxation time approximation [31]. This heuristic $MP$ parameter accounts for (i) the effect of a finite mean free path in real materials relative to the ballistic transport assumed in our simulations, (ii) the difference in dimensionality between our nanoscale models and the larger experimental thin-film system , (iii) the approximate correction of the band gap we employed, and (iv) the compensation for the partially arbitrary choice of an auxiliary function in the model (which will be described later in the text).
@@
We relate the electronic conductances of our NW models to experimental electric conductivity data available in the literature. In Figure 13, panel b, it can be observed that the BT 155v systems exhibit the highest conductivity among the systems we have examined. We associate this model with an hypothetical real ScN system characterized by a strong predominance of N-site vacancies occurring in contiguous positions. The experimental ScN sample from Ref. 48 exhibits some of the highest electric conductivity values ever reported for ScN thin films.
@@
We assume that a strong presence of structural vacancies at N-sites, associated in contiguous positions, is at the origin of the high experimental conductivity. It should be noted that the association of vacancies at N-sites in contiguous positions is not thermodynamically favored. We treat surface defects with the idea to describe the deposition process, and in order to quantify the energy involved, we performed calculations of the formation enthalpies of isolated surface vacancies at N-sites in a ScN slab structure, as well as the formation of associated defects  [^2] We calculated the formation energies of associated defects in 6-atomic-layer slab structures by means of DFT as described in section II.3, using a 3×3 k-point grid and exploiting space-group symmetries where possible. The periodic unit used for these calculations had in-plane dimensions of 3×3 conventional simple cubic cells. Atomic layer 1 was kept fixed during the relaxation calculations, and defects were inserted in layers 4 and 6 at the in-plane center of the periodic unit, in axial positions perpendicular to the slab.. The formation enthalpy of two contiguous vacancies results higher by 8.92% compared to that of two isolated vacancies (5.44 eV per single isolated vacancy). However, the formation of a single vacancy at N-site is itself thermodynamically unfavorable, and the first studies on the stability of nitrogen vacancy defects [56] argued that these are not energetically favored and were therefore expected not to be significantly present in ScN lattices. More recent experimental reports have clearly shown, instead, that nitrogen vacancies are prevalent, especially under scandium-rich growth conditions [5, 65]. Although these conditions are not easily achieved with sputtering techniques, this suggests the existence of complex mechanisms for the stabilization of nitrogen vacancies that are not yet fully understood. For instance, the formation of void regions in the bulk structure of ScN has been shown to facilitate a higher concentration of n-type impurities in their vicinity [38], and other theoretical studies have demonstrated a negative formation energy for surface N-site vacancies under Sc-rich conditions [66].
@@
These reasons moved to our assumption and therefore we derived a single model parameter $MP$, which we used for all the systems presented in Figure 13, by fitting the BT 155v conductivity to the experimental data reported in Ref. 48, corresponding to the sample referred to in that work as ScN (here we denote it as Sample I). By intersecting a model-internal theoretical auxiliary function, which we report in Figure 13, panel b (155v aux.), we derived the chemical potentials of that sample at various temperatures from the experimental Seebeck coefficients reported in the same work [48]. As auxiliary function, we used the theoretical Seebeck function of each NW model itself, calculated at a temperature of 450 K (this choice is partially arbitrary and is compensated by the derivation of a suitable $MP$, as previously discussed). We obtained the best possible fit of the model by associating the auxiliary function of the 155 model with the experimental sample labeled ScN-T in Ref. 48 (here denoted Sample II), the auxiliary function of the BT 155v model with the experimental sample labeled ScN in Ref. 48 (Sample I, used to fit $MP$), and the auxiliary function of BT 155O+sf with the experimental sample examined in Ref. 12 (here denoted Sample III). In Figure 13, we report these three auxiliary functions and the calculated conductivities comparing with experimental values from Refs. 12 and 48.
@@
We chose to compare the two models of defected structures noted before that present the most interesting features when comparing their Seebeck function profiles with that of the pristine 155 model, namely: BT 155v and BT 155O+sf. The compared experimental samples (Refs. 12 and 48), deriving from different experimental procedures, differ widely from each other in their characteristics, presenting a wide variability in Seebeck coefficients, electric conductivities, and their variance with temperature. We notice a good agreement between the experimental conductivity values measured for the ScN-T sample in Ref. 48 (Sample II) and the theoretical functions calculated using the NW model 155O+sf (we mention a slightly lower, yet still good, agreement when using the auxiliary function of the 155O model, not reported in figure).
@@
We can therefore describe, within a coherent framework, an interaction between two independent effects that determine the different thermoelectric behaviors of the ScN and ScN-T samples in Ref. 48 (Sample I and II). The different experimental procedures lead, on one hand, to a reduction in v+v defects in ScN-T compared to ScN. This reduction underlies the strong increase in the absolute value of the Seebeck coefficient observed in Ref. 48 and also contributes to the decrease in electrical conductivity. On the other hand, an additional detrimental effect on electronic transport arises from a greater presence of O+sf defects.
@@
Further complexity is detected when analyzing the experimental sample from Ref. 12 (Sample III). This sample also exhibits high conductivity values, so it is reasonable to expect its experimental values to match the highest conductivity function, namely the one calculated using the 155v model. However, in order to achieve good agreement, a different auxiliary function must be used in this case. The best match is obtained when using an auxiliary function calculated from the BT 155O+sf system. The choice of a different auxiliary function accounts for structural effects that may influence the Seebeck coefficient but have little impact on the conductivity. These considerations suggest that different defected structures can coexist within the same sample, acting independently on the electric conductivity and the Seebeck coefficient, with possible contrasting effects.
@@
At this point, it should be noted that while the 155v model has a physically unique structure determined by a structural optimization calculation, the structure used for the 155O+sf model involves a greater degree of arbitrariness in the way the planar mismatch was introduced as a stacking fault. As previously discussed, the thermoelectric properties of the 155O+sf model are highly structure-dependent. From an analytical perspective, the shape of its auxiliary function is well suited to fit the model. However, to achieve better numerical agreement, one would need to use the Seebeck function calculated at 600 K as auxiliary function (instead of the one at 450 K, which we use). We instead choose to preserve the internal consistency of the model by calculating all auxiliary functions in the same way. As a result, this set of points yields a lower agreement between theory and experiment compared to the others. However, the reasons for this outcome are clear, as the 155O+sf defect category inherently includes a greater amount of structural variability.
To summarize, based on the results of our theoretical calculations, we derived a heuristic model that relates our simulated atomistic systems to the experimental conductivities reported in Ref. 48 and 12. We used the data of Sample I (ScN, Ref. 48) to fit the model. We obtained good agreement with the experimental conductivities of Sample II (ScN-T, Ref. 48). A slightly weaker agreement is found with Sample III (ScN, Ref. 12).
@@
Figure 13: Panel a: Auxiliary functions employed to relate the experimental Seebeck coefficients to theoretical values of chemical potential. The figure reports the functions of pristine and two different BT defected (1x)5x5+s ScN NW structures (v+v and O+sf). These functions are calculated as the Seebeck functions at 450 K. The figure also reports, for a comparison, the Seebeck function calculated at 600 K for the BT 155O+sf structure. Panel b: The lines represent calculated electric conductivities as functions of the chemical potential for the same three model systems, at different temperatures. Color intensities of the lines increase with temperature from 300 to 780 K. The points represent, at their calculated chemical potentials, experimental values of electric conductivities: Sample I: ScN from Ref. 48, Sample II: ScN-T from Ref. 48 and Sample III: ScN from Ref. 12.
@@ IV Conclusions
We wondered what the microscopic origins of the experimental variability in thermoelectric response of ScN thin films might be. Knowing that the thermoelectric properties of ScN are strongly influenced by stoichiometric variability, structural imperfections, and chemical impurities, we decided to model some of these structural features to understand how they act in modifying the electronic transport, and consequently the thermoelectric response.
We employed a theoretical framework suitable for studying in detail these atomistic effects on quantum transport. We approached the problem by classifying the possible structural defects and impurities based on their symmetry and chemical nature, and we systematically analyzed those most likely to occur under the experimental conditions typically used.
From the point of view of defect symmetry, we considered isolated point defects, multiple contiguous point defects, and point defects associated with planar mismatches (stacking faults), while from a chemical perspective we studied the presence of oxygen or structural vacancies replacing nitrogen atoms in the lattice. We used the Landauer approach to determine the effect of each of these different microscopic structures on the emergent thermoelectric properties of the whole material.
Thermoelectric conversion is a complex phenomenon and electronic transport is certainly influenced, through electron-phonon scattering, by extended defects such as nanovoids, grain orientation, substrate-induced strain. Therefore our theoretical approach cannot account for the full extent of the experimental variability. Nevertheless, electronic transport is also fundamentally related to the individual quantum states and to the microscopic regions where the single electrons flow.
We identified two notable categories of structures whose presence has a considerable impact on the Seebeck coefficient and electrical conductivity at the microscopic level and which, if significantly present in the material, could partially account for the experimentally observed behavior.
Specifically, we refer first to the v-v structures, where two or more structural vacancies replace nitrogen atoms at adjacent conventional cells (i.e., N-site vacancies separated by only one Sc atom), which can significantly reduce the absolute value of the Seebeck coefficient while enhancing the electrical conductivity. Second, we refer to the O+sf structures, where an oxygen atom replaces a nitrogen lattice atom, and a stacking fault is immediately adjacent, involving a displacement of the atomic layer following the one where the substitution occurs. This second category of structures has potentially opposite effects with respect to the firts one, detrimenting the conductivity while enhanching the Seebeck coefficient.
Since n-type defects tend to accumulate in the vicinity of nanovoid regions in the ScN lattice, the proximity of these structures could allow v-v and O+sf formations to reach a concentration sufficient to produce an appreciable effect on the thermoelectric response of experimental samples, offering a rationale for part of the observed experimental variability.